OpenStudy (anonymous):

\[tan^{-1}0\] \[tan(...)=0\] that would be \[\frac \pi 2\] so \[tan^{-1}0=\frac \pi 2\]

OpenStudy (anonymous):

@satellite73 Am I right?

OpenStudy (anonymous):

i mean yes!!

OpenStudy (anonymous):

sorry i am tired, i thought you wanted \(\tan^{-1}(1)\) you are right

OpenStudy (anonymous):

\[tan^{-1}\sqrt 3=\frac\pi 6\]

OpenStudy (anonymous):

now that one is wrong

OpenStudy (anonymous):

noooooo....

OpenStudy (anonymous):

\[\tan(\frac{\pi}{3})=\sqrt3\]

OpenStudy (anonymous):

oh it's \[\frac \pi 3\]

OpenStudy (anonymous):

so \[\tan^{-1}(\sqrt3)=\frac{\pi}{3}\]

OpenStudy (anonymous):

yeah tan is \[\frac yx\] yep makes sense haha

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

awesome!